OpenStudy (anonymous):

(2a^2 - a + 2) ----------- - (a) -1 pls help (a-1)

OpenStudy (anonymous):

What do you need help with? Simplifying, or...?

OpenStudy (anonymous):

simplifying this subtration of rational expression

OpenStudy (anonymous):

\[=\frac{(2a^2 - a + 2 -a^2+a-a+1)}{(a-1)}\] \[=\frac{(3a^2-a+3)}{(a-1)}\]

OpenStudy (anonymous):

ei wrong answer